# Do Visual-Language Grid Maps Capture Latent Semantics?

### v2

@@ I Introduction
    We make our evaluation software publicly available [^1] The source code will be made publicly available upon acceptance of this work., enabling evaluation of the progress in this new field of research.
@@ IV-A Evaluating queryability
    Each map $m\in\mathcal{M}$ is queried with each query $q$ in the set of queries $\mathcal{Q}$. The query result of a single query is a binary segmentation mask of the map, $\hat{\textbf{y}}_{q}=\{\hat{y}_{1},\dots,\hat{y}_{N}\}$, consisting of $N$ voxel masks $\hat{y}\in\{true,false\}$. Combined, the query results form the set of voxel-based predictions $\hat{Y}_{V}=\{\hat{\textbf{y}}_{q}|q\in\mathcal{Q}\}$.
@@
    For each query, the true mask $\textbf{y}_{q}$ is created from the ground truth map $m_{g}$ such, that the voxels answering query $q$ are $true$, others $false$. The true masks form a set of true labels $Y_{V}=\{\textbf{y}_{q}|q\in\mathcal{Q}\}$.
@@ IV-B1 Intra-map distinctness
    The average absolute deviation $d_{l}^{i}$ is calculated for $\mathcal{E}_{l}$, and $d_{\mu}^{i}$ for $\mathcal{E}_{i}$, for each tuple $t\in\mathcal{T}$. The average absolute deviation $d_{a}$ measures the statistical dispersion of a set, and given a set of points =$\mathcal{E}=\{e_{1},...,e_{n}\}$ it is defined as
@@ V-A Data set
The data set used in the study is the Matterport3D data set [31], which consists of houses captured with an RGBD camera and ground truth semantic segmentation. The data set was used in the original work of VLMaps, in which they selected a sample of 10 sequences from the data set. In this work, we use the same sequences and the same set of 40 labels [31] forming the set $\mathcal{L}$; their names are also used in the following experiments as the query set $\mathcal{Q}$ to leverage the semantic ground truth provided with the dataset [^2] Supplementary material including the list of labels will be available in arXiv and link provided here upon acceptance of this work..
@@ V-C Querying the maps
VLMaps proposes a method for performing open-vocabulary queries, which we will refer to as VLMaps querying. When the map is queried with a query $q$, the embeddings on the map are compared using cosine similarity to two embeddings, embedding of $q$ and embedding of string “other”. The binary classification mask is $true$ for voxels where the most similar query is $q$; otherwise, it is $false$. The binary mask is then post-processed by a sequence of binary closing, gaussian blurring, thresholding, and binary dilation to encompass whole object instances.
@@ V-D Resolution analysis
For real-world applications, where the robot hardware capabilities might be limited, the memory footprint of the map can be crucial. Small cell sizes can lead to prohibitively large maps, and the size of a grid map is a function of the resolution of the map, as shown in Figure 2. OpenScene has smaller map sizes due to downsampling the images. To investigate the effects of cell size on the map quality, all of the experiments were performed with a set of resolutions $r\in[0.02,0.05,0.1,0.2]$ m. The resolutions proposed in the original works are $0.02$m and $0.05$m in OpenScene and VLMaps, respectively. If not explicitly stated otherwise, we show the results in a resolution of $0.02$m for fair comparison.

### v3

@@ I Introduction
    We make our evaluation software publicly available [^1] https://aalto-intelligent-robotics.github.io/latent_semantics/, enabling evaluation of the progress in this new field of research.
@@ IV-A Evaluating queryability
[…]
@@
    For each query, the true mask $\textbf{y}_{q}$ is created from the ground truth map $m_{g}$ such, that the voxels answering query $q$ are truetrue, others falsefalse. The true masks form a set of true labels $Y_{V}=\{\textbf{y}_{q}|q\in\mathcal{Q}\}$.
@@ IV-B1 Intra-map distinctness
    The average absolute deviation $d_{l}^{i}$ is calculated for $\mathcal{E}_{l}$, and $d_{\mu}^{i}$ for $\mathcal{E}_{i}$, for each tuple $t\in\mathcal{T}$. The average absolute deviation $d_{a}$ measures the statistical dispersion of a set, and given a set of points $\mathcal{E}=\{e_{1},...,e_{n}\}$ it is defined as
@@ V-A Data set
The data set used in the study is the Matterport3D data set [31], which consists of houses captured with an RGBD camera and ground truth semantic segmentation. The data set was used in the original work of VLMaps, in which they selected a sample of 10 sequences from the data set. In this work, we use the same sequences and the same set of 40 labels [31] forming the set $\mathcal{L}$; their names are also used in the following experiments as the query set $\mathcal{Q}$ to leverage the semantic ground truth provided with the dataset [^2] https://aalto-intelligent-robotics.github.io/latent_semantics/.
@@ V-C Querying the maps
VLMaps proposes a method for performing open-vocabulary queries, which we will refer to as VLMaps querying. When the map is queried with a query $q$, the embeddings on the map are compared using cosine similarity to two embeddings, embedding of $q$ and embedding of string “other”. The binary classification mask is truetrue for voxels where the most similar query is $q$; otherwise, it is falsefalse. The binary mask is then post-processed by a sequence of binary closing, gaussian blurring, thresholding, and binary dilation to encompass whole object instances.
@@ V-D Resolution analysis
For real-world applications, where the robot hardware capabilities might be limited, the memory footprint of the map can be crucial. Small cell sizes can lead to prohibitively large maps, and the size of a grid map is a function of the resolution of the map, as shown in Figure 2. OpenScene has smaller map sizes due to downsampling the images. To investigate the effects of cell size on the map quality, all of the experiments were performed with a set of resolutions $r\in[0.02,0.05,0.1,0.2]$ m. The resolutions proposed in the original works are 0.020.02m and 0.050.05m in OpenScene and VLMaps, respectively. If not explicitly stated otherwise, we show the results in a resolution of 0.020.02m for fair comparison.